6. A 0.3 kg rubber ball is dropped from the window of a building. It strikes the sidewalk below at 30 m/s and rebounds up at 20 m/s. The magnitude of the impulse due to the collision with the sidewalk is:


1
Expert's answer
2021-08-16T09:43:26-0400

The magnitude of the impulse due to the collision of a rubber ball with the sidewalk

"\Delta p=m|{\bf v_2-v_1}|=m(v_2-(-v_1))\\=0.3\:\rm kg*(20+30)\: m/s=15\: kg*m/s"